question 9 and 10Question 9 (10 points) Kramerica Industires plans to introduce a new product to the market. Last week, Kramerica hired a mark
Answer #1

Calculation of cash flows in year 20:

Incremental Sales = 9,000,000*80% = $7,200,000

Less: variable costs = $3,400,000

Fixed costs = $700,000

Depreciation (26 million – 4 million)/20 = $1,100,000

Income before tax = $2,000,000

Less: Tax @40% = $800,000

Income after Tax = $1,200,000

Add: Depreciation = $1,100,000

Cash flow from operations = $2,300,000

Add: Salvage value of Equipment = $4,000,000

Add: Recovery of working capital = $300,000

Cash flow in year 20 = $6,600,000

10.The answer is

The property on which the plant will be built has bought in 1950 for $10,000

This is a sunk cost already incurred and hence not relevant